How do you solve this equation using the elimination method?
a+7b=7
a+2b=-3

Answer by Ave(106)

a+7b=7
a+2b=-3 subtract
5b=10 divide by 5
b=2
now plug in 2 for b in either equation and solve for a
a+2(2)=-3
a+4=-3
a=-7
---
a=-7, b=2